The short run total cost function is SRTC(Q) = $1,200,000 + 200Q

What is the lowest price which the firm can accept for one unit of Q?
Answer #1

The average variable cost = d/dQ (SRTC) = d/dQ (1,200,000 +200Q) = 200.

So, the minimum acceptable price for the firm is $200 as this is the average variable cost.